OpenStudy (anonymous):

What is the y-coordinate of the vertex of y = |–3x + 6| – 2 ? i got 0 but i'm not 100% sure

OpenStudy (anonymous):

It will be -2 since the lowest value |-3x+6| reaches is zero, since absolute values are never negative. And the vertex of such an absolute value function occurs on the axis of symmetry, which is where the absolute value part is zero.

OpenStudy (anonymous):

y=-3x+6-2, for x <2 and y =3x-6-2 for x>2 use this to draw the graph and get the vertex

OpenStudy (anonymous):

it can be y=(-3x+6)-2 or y=3x-6-2 then y=-3x+4 or y=3x-8

OpenStudy (anonymous):

okay, well its not -2... and @sriramkumar now what? solve it or is there another step... a little confused now

OpenStudy (anonymous):

now, the answer depends on the x-co ordiante you have @jeyde205 when u have an x value, subsitute in the above eqns i have deduced and you can get the y values...

OpenStudy (anonymous):

okayy thanks:)

OpenStudy (anonymous):

 @jeyde205 solve the two equations , i had mentioned and the value of x and y you get after solving will be the vertex